The number of calls received by an office on Monday morning between 8:00 AM and 9:00 AM has a mean of 2. Calculate the probability of getting at least 2 calls between eight and nine in the morning. Round your answer to four decimal places.
Answer #1

X ~ Poi ( )

Where = 2

Poisson probability distribution is

P(X) = e- * X / X!

So,

P(X >= 2) = 1 - P(X <= 1)

= 1 - [ P(X = 0) + P(X = 1) ]

= 1 - [ e-2 + e-2 * 2]

= 0.5940